Magic Realm, Inc., has developed a new fantasy board game. The company sold 44,000 games last year at a selling price of $64 per game. Fixed expenses associated with the game total $792,000 per year, and variable expenses are $44 per game. Production of the game is entrusted to a printing contractor. Variable expenses consist mostly of payments to this contractor.

Required:

1-a. Prepare a contribution format income statement for the game last year.

1-b. Compute the degree of operating leverage.

2. Management is confident that the company can sell 53,240 games next year (an increase of 9,240 games, or 21%, over last year). Given this assumption:

a. What is the expected percentage increase in net operating income for next year?

b. What is the expected amount of net operating income for next year? (Do not prepare an income statement; use the degree of operating leverage to compute your answer.)
Answer #1

1a) Contribution margin income statement

Sales (44000*64) 2816000
Variable cost 1936000
Contribution margin 880000
Fixed cost 792000
Net operating income 88000

1b) Degree of operating leverage = 880000/88000 = 10

2a) Net operating income increase by = 21*10 = 210%

2b) Expected net income = 88000*3.1 = 272800